OpenStudy (anthony1914):

Simplify (8x2 − 1 + 2x3) − (7x3 − 3x2 + 1).

OpenStudy (kendricklamar2014):

@anthony1914 What have you done to solve this problem? Can you show me the steps you took to answer this?

OpenStudy (skullpatrol):

Any ideas? :-)

OpenStudy (skullpatrol):

Subtracting polynomials is like subtracting real numbers. Add the opposite of each term of the polynomial and simplify.

OpenStudy (jjcoolja12):

Let's simplify step-by-step. 8x2−1+2x3−(7x3−3x2+1) Distribute the Negative Sign: =8x2−1+2x3+−1(7x3−3x2+1) =8x2+−1+2x3+−1(7x3)+−1(−3x2)+(−1)(1) =8x2+−1+2x3+−7x3+3x2+−1 Combine Like Terms: =8x2+−1+2x3+−7x3+3x2+−1 =(2x3+−7x3)+(8x2+3x2)+(−1+−1) =−5x3+11x2+−2 Answer: =−5x3+11x2−2